OpenStudy (anonymous):

I need this solved out in steps.2(x + 6) − 10 = 12 + 4(x − 1)

OpenStudy (aravindg):

first use distributive property ...Then add like terms

OpenStudy (anonymous):

2x + 12 - 10 = 12 + 4x - 4 2x + 2 = 8 + 4x 2x + 2 - 8 - 4x = 0 -2x - 8 = 0 -2x = 8 x = 14 I am not sure, but this is likely to be the answer...

OpenStudy (texaschic101):

2(x + 6) -10 = 12 + 4(x - 1) -- distribute through the parenthesis 2x + 12 - 10 = 12 + 4x - 4 -- combine like terms 2x + 2 = 8 + 4x -- subtract 8 from both sides 2x + 2 - 8 = 8 - 8 + 4x -- combine like terms 2x - 6 = 4x -- subtract 2x from both sides 2x - 2x - 6 = 4x - 2x -- combine like terms -6 = 2x -- divide by 2 to get x by itself -6/2 = 2x/2 -3 = x check... 2(x + 6) - 10 = 12 + 4(x - 1) 2(-3 + 6) - 10 = 12 + 4(-3-1) 2(3) - 10 = 12 + 4(-4) 6 - 10 = 12 - 16 - 4 = - 4 (correct) ANSWER : x = - 3 I tried to do it step by step....if you have any questions, please ask them :)

OpenStudy (anonymous):

sorry i meant -4 not 14